# A Survey on Data-Centric Recommender Systems

### v1

@@ 4.2.3 Dataset Distillation/Condensation
Dataset distillation or dataset condensation techniques Yu et al. (2023) aim to synthesize data points with the goal of condensing the comprehensive knowledge from the entire dataset into a small, synthetic data summary. This process retains the essence of the data, enabling models to be trained more efficiently. Recently, some works Sachdeva et al. (2022); Wu et al. (2023) observed that dataset distillation has a strong data denoising effect in recommendation. For example, DISTILL-CF Sachdeva et al. (2022) leverages dataset meta-learning to synthetic user-item interactions. Remarkably, models trained on the condensed dataset synthesized by DISTILL-CF have demonstrated improved performance compared to those trained on the full, original dataset.
@@ 4.2.4 Discussion
Normally, data collected from real-world scenarios are frequently contaminated with noise – stemming from system bugs or user mistakes. However, obtaining labels for this noise is often impractical or even impossible, due to the lack of expert knowledge required for identifying noise, the high costs associated with manual labeling, or the dynamic nature of some noise which makes it hard to give fixed labels. In the absence of the ground truth, it is difficult to determine whether the denoising method achieves the optimal situation – no over-denoising or under-denoising. Taking into account this limitation, it may be preferable to synthesize a noise-free dataset via dataset distillation/condensation rather than attempting to adjust or filter the existing dataset through reweighting-based or selection-based methods.
@@ 4.3.1 User Preference Alignment
User preferences may shift due to a variety of reasons, such as temporal changes Zafari et al. (2019); Wangwatcharakul and Wongthanavasu (2021); Ding et al. (2022), locational moves Yin et al. (2016), or alterations in personal and environmental conditions Zheng et al. (2021); He et al. (2022); Wang et al. (2023a). Existing methods are designed to track and adjust to these changes, thereby maintaining alignment with the ever-evolving user preferences. For example, Aspect-MF Zafari et al. (2019) analyzes and models temporal preference drifts using a component-based factorized latent approach. MTUPD Wangwatcharakul and Wongthanavasu (2021) adopts a forgetting curve function to calculate the correlations of user preferences across time. ST-LDA Yin et al. (2016) learns region-dependent personal interests and crowd preferences to align locational preference drifts. Wang et al. Wang et al. (2023a) review user preference shifts across environments from a causal perspective and inspect the underlying causal relations through causal graphs. Based on the causal relations, they further propose the CDR framework, which adopts a temporal variational autoencoder to capture the preference shifts.
@@ 4.3.3 Discussion
Traditional evaluation settings in RS may not be appropriate for assessing debiasing methods because they typically entail that the distribution of a test set is representative of the distribution in the training set (independent and identically distributed evaluation settings). Therefore, in this part, we discuss two out-of-distributed evaluation settings to assess debiasing methods:
  • Temporal split setting: Temporal split setting slices the historical interactions into the training, validation, and test sets according to the timestamps Zhang et al. (2022a, 2023b). In that case, any shift in user preferences or item popularity over time is appropriately represented and accounted for during the evaluation.
@@ 5 Future Directions
5.1 Data-Centric RS with Multimodal Data
Multimodal data refers to data that consists of multiple modalities or types of information, such as text, images, audio, video, or any combination thereof. Traditionally, RS have primarily relied on user-item interaction data, such as ratings or click-through data, to generate recommendations. However, by incorporating multimodal data, RS can capture richer and more diverse user preferences and item characteristics, leading to more personalized and relevant recommendations Truong and Lauw (2019). However, the data issues mentioned before (i.e., incomplete data, noisy data, and biased data) also exist in multimodal data, and they can pose additional challenges in the context of multimodal RS:
@@ 5.1 Data-Centric RS with Multimodal Data
  • Imbalance: Multimodal datasets may exhibit imbalances in the distribution of different modalities. For example, there may be a larger number of text samples compared to images or audio samples. Modality imbalance can affect the performance and generalization of recommendation models trained on such data.
@@
5.2 Data-Centric RS with Large Language Models
With the emergence of large language models (LLMs) in Natural Language Processing, there has been a growing interest in harnessing the power of these models to enhance recommender systems. In Data-Centric RS, LLMs can serve as:
@@ 5.2 Data-Centric RS with Large Language Models
  • Data processors: As mentioned before, given the extensive knowledge base and powerful reasoning capabilities of LLMs, some recent work has attempted to augment data with LLMs, for example, through carefully designed prompts, LLMRec Wei et al. (2023) employs three simple yet effective LLM-based data augmentation strategies to augment implicit feedback, user profiles, and item attributes, respectively. Moving forward, it’s crucial to investigate the capability of LLMs in managing tasks such as data denoising and data debiasing. This could pave the way for LLMs to harmonize Data-Centric RS effectively.
5.3 Automatic Data-Centric RS
Automatic machine learning (AutoML) He et al. (2021) refers to the process of automating the end-to-end process of applying machine learning to real-world problems, which typically involves automating a variety of tasks that are part of the machine learning workflow. In the context of Data-Centric RS, these tasks encompass data augmentation, data denoising, and data debiasing. Consequently, AutoML can automatically streamline and enhance the efficiency of these tasks, enabling more accurate recommendations, which is of great significance in application.
5.4 Transparent Data-Centric RS
Transparent Data-Centric RS refer to Data-Centric RS that not only offer enhanced data for model training but also provide insights into how and why particular enhancements are made, thereby allowing users and developers to understand the underlying decision-making processes. Research in transparent Data-Centric RS is tackling complex challenges, such as the trade-off between transparency and complexity, ensuring user privacy while providing explanations, and developing standards for explainability and interpretability.
@@ 6 Conclusions
In this survey, we conclude a comprehensive literature review of Data-Centric RS. We systematically analyze three critical issues inherent in recommendation data and subsequently categorize existing works in accordance with their focus on these issues. Additionally, we point out a range of prospective research directions that advance Data-Centric RS. We hope this survey can help the readers efficiently grasp a broad picture of this field, and equip them with the techniques and further research ideas to build recommender systems.

### v4

@@ 4.2.3 Dataset Distillation/Condensation
Dataset distillation or dataset condensation techniques Yu et al. (2023) aim to synthesize data points with the goal of condensing the comprehensive knowledge from the entire dataset into a small, synthetic data summary. This process retains the essence of the data, enabling models to be trained more efficiently. Recently, some works Sachdeva et al. (2022); Wu et al. (2023) observe that dataset distillation has a strong data denoising effect in recommendation. For example, DISTILL-CF Sachdeva et al. (2022) applies dataset meta-learning to synthetic user-item interactions. Remarkably, models trained on the condensed dataset synthesized by DISTILL-CF have demonstrated improved performance compared to those trained on the full, original dataset.
@@ 4.2.4 Discussion
Normally, data collected from real-world scenarios are frequently contaminated with noise stemming from system bugs or user mistakes. However, obtaining labels for this noise is often impractical or even impossible, due to the lack of expert knowledge required for identifying noise, the high costs associated with manual labeling, or the dynamic nature of some noise which makes it hard to give fixed labels. In the absence of the ground truth, it is difficult to determine whether the denoising method achieves the optimal situation – no over-denoising or under-denoising. Taking into account this limitation, it may be preferable to synthesize a noise-free dataset via dataset distillation/condensation rather than attempting to adjust or filter the existing dataset through reweighting-based or selection-based methods.
@@ 4.3.1 User Preference Alignment
User preferences may shift due to a variety of reasons, such as temporal changes Zafari et al. (2019); Wangwatcharakul and Wongthanavasu (2021); Ding et al. (2022), locational moves Yin et al. (2016), or alterations in personal and environmental conditions Zheng et al. (2021); He et al. (2022); Wang et al. (2023a). Existing methods are designed to track and adjust to these changes, thereby maintaining alignment with the ever-evolving user preferences. For example, Aspect-MF Zafari et al. (2019) analyzes and models temporal preference drifts using a component-based factorized latent approach. MTUPD Wangwatcharakul and Wongthanavasu (2021) adopts a forgetting curve function to calculate the correlations of user preferences across time. ST-LDA Yin et al. (2016) learns region-dependent personal interests and crowd preferences to align locational preference drifts. Wang et al. Wang et al. (2023a) review user preference shifts across environments from a causal perspective and inspect the underlying causal relations through causal graphs. Based on the causal relations, they further propose the CDR framework, which adopts a temporal variational autoencoder to capture preference shifts.
@@ 4.3.3 Discussion
Traditional evaluation settings in RSs may not be appropriate for assessing debiasing methods because they typically entail that the distribution of a test set is representative of the distribution in the training set (independent and identically distributed evaluation settings). Therefore, in this part, we discuss two out-of-distribution evaluation settings to assess debiasing methods:
  • Temporal split setting: Temporal split setting slices the historical interactions into the training, validation, and test sets according to the timestamps Zhang et al. (2022a, 2023b). In this case, any shift in user preferences or item popularity over time is appropriately represented and accounted for during the evaluation.
@@ 5 Future Directions
5.1 Data-Centric RSs with Multimodal Data
Multimodal data refers to data that consists of multiple modalities or types of information, such as text, images, audio, video, or any combination thereof. Traditionally, RSs have primarily relied on user-item interaction data, such as ratings or click-through data, to generate recommendations. By incorporating multimodal data, RSs can capture richer and more diverse user preferences and item characteristics, leading to more personalized and relevant recommendations Truong and Lauw (2019). However, the data issues mentioned before (i.e., incomplete data, noisy data, and biased data) also exist in multimodal data, and they can pose additional challenges in the context of multimodal RSs:
@@ 5.1 Data-Centric RSs with Multimodal Data
  • Imbalance: Multimodal datasets may exhibit imbalances in the distributions of different modalities. For example, there may be a larger number of text samples compared to images or audio samples. Modality imbalance can affect the performance and generalization of recommendation models trained on such data.
@@
5.2 Data-Centric RSs with LLMs
With the emergence of large language models (LLMs) in natural language processing, there has been a growing interest in harnessing the power of these models to enhance recommender systems. In Data-Centric RSs, LLMs can serve as:
@@ 5.2 Data-Centric RSs with LLMs
  • Data processors: As mentioned before, given the extensive knowledge base and powerful reasoning capabilities of LLMs, some recent work has attempted to augment data with LLMs, for example, through carefully designed prompts, LLMRec Wei et al. (2023) employs three simple yet effective LLM-based data augmentation strategies to augment implicit feedback, user profiles, and item attributes, respectively. Moving forward, it’s crucial to investigate the capability of LLMs in managing tasks such as data denoising and data debiasing. This could pave the way for LLMs to harmonize Data-Centric RSs effectively.
5.3 Automatic Data-Centric RSs
Automatic machine learning (AutoML) He et al. (2021) refers to the process of automating the end-to-end process of applying machine learning to real-world problems, which typically involves automating a variety of tasks that are part of the machine learning workflow. In the context of Data-Centric RSs, these tasks encompass data augmentation, data denoising, and data debiasing. Consequently, AutoML can automatically streamline and enhance the efficiency of these tasks, enabling more accurate recommendations, which is of great significance in practice.
5.4 Transparent Data-Centric RSs
Transparent Data-Centric RSs refer to Data-Centric RSs that not only offer enhanced data for model training but also provide insights into how and why particular enhancements are made, thereby allowing users and developers to understand the underlying decision-making processes. Research in transparent Data-Centric RSs is tackling complex challenges, such as the trade-off between transparency and complexity, ensuring user privacy while providing explanations, and developing standards for explainability and interpretability.
@@ 6 Conclusions
In this survey, we presented a comprehensive literature review of Data-Centric RSs. We systematically analyzed three critical issues inherent in recommendation data and subsequently categorized existing works in accordance with their focus on these issues. Additionally, we point out a range of prospective research directions to advance Data-Centric RSs. We expect that this survey can help readers easily grasp the big picture of this emerging field and equip them with the basic techniques and valuable future research ideas.